OpenStudy (anonymous):

solve: dy/dt = 0.12y, where y(0) = 3

OpenStudy (unklerhaukus):

separate the variables, and integrate

OpenStudy (anonymous):

Separate the variables?

sam (.sam.):

\[\frac{dy}{dt} = 0.12y \\ \\ \frac{1}{y}dy=0.12dt\]

sam (.sam.):

Integrate both sides

OpenStudy (anonymous):

What is the reasoning behind this? If I were looking at this on a test, how would I know to separate the variables and integrate?

sam (.sam.):

You basically want to separate the y's with the dy and t's with dt

sam (.sam.):

Then you will able to integrate this

OpenStudy (anonymous):

And if I do this, I end up with ln(y)=0.12t, but I don't know what to do from there.

terenzreignz (terenzreignz):

Please don't forget the constant of integration... \[\Large \ln(y)= 0.12t + C_o\]

terenzreignz (terenzreignz):

And from here, you can take it further...

terenzreignz (terenzreignz):

You generally want the answer to a differential equation (that's what this kind of problem is called) as a function, generally still, a function of t. What you can do is raise e to both sides of the equation... \[\LARGE e^{\ln(y)}=e^{0.12t + C_o}\]\[\LARGE y = e^{C_o}e^{0.12t}\] And to simplify things, since Co was an arbitrary constant, then \(\Large e^{C_o}\) is also just some constant, so... \[\huge y = \color{blue}Ce^{0.12t}\]

OpenStudy (anonymous):

Amazing, thank you very much.

terenzreignz (terenzreignz):

I just polished the edges :)

terenzreignz (terenzreignz):

And if you're still here, allow me to present a problem where separation of variables is a more pressing need... oh, and if any of the following is confusing, just ignore them... the last thing we want is to overcomplicate things :D \[\Large \frac{dy}{dt}=2t^2y -2t^2+3ty-3t\] does it look menacing? :)

terenzreignz (terenzreignz):

Upon closer inspection, you'll find that "menacing" expression to the right may actually be factored... \[\Large \frac{dy}{dt}= (y-1)(2t^2 +3t)\]

terenzreignz (terenzreignz):

So, at this point, you may treat the dy and the dt as variables in their own right, which may be multiplied, divided, cancelled, etc. In particular, multiply both sides of the equation by "dt" \[\Large \cancel{dt} \cdot \frac{dy}{\cancel{dt}}=(y-1)(2t^2+3t)dt\]

terenzreignz (terenzreignz):

And divide both sides of the equation by (y-1) \[\Large \frac1{y-1}dy = (2t^2+3t)dt\]

terenzreignz (terenzreignz):

And now, what was once a rather "menacing" looking differential equation is ripe for integration. \[\Large \int\frac1{y-1}dy =\int (2t^2+3t)dt\]

terenzreignz (terenzreignz):

Skipping all the nitty-gritty details of integration, and bringing all constants of integration to the right-hand-side, we arrive at \[\Large \ln(y-1) = \frac23t^3 + \frac32t^2+C_o\]

terenzreignz (terenzreignz):

Using a similar trick, we raise e to both sides of the equation to get rid of the "ln" at the left side, such that this expression may be rearranged as a function of t. \[\LARGE e^{\ln(y-1)}=e^{\frac23t^3+\frac32t^2+C_o}\] \[\LARGE y-1 = e^{C_o}\times e^{\frac23t^3 + \frac32t^2}\] Finally \[\LARGE y = Ce^{\frac23t^3 + \frac32 t^2}+1\]

OpenStudy (anonymous):

In the first part, how did e^C become simply C?

terenzreignz (terenzreignz):

Because Co was just some arbitrary constant, e^Co would also just be an arbitrary constant, so I just replaced e^Co with C, for simplicity